40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 20 19:09:48 2015, ericq, Update, ASC, QPD frontend code unified 
    Reply  Mon Feb 23 08:06:59 2015, rana, Update, ASC, QPD frontend code unified 
    Reply  Thu Mar 5 20:44:30 2015, Jenne, Update, SUS, damprestore script updated 
    Reply  Mon Mar 9 18:10:59 2015, Jenne, Update, ASC, Oplev filters re-copied 
    Reply  Sat Mar 14 00:12:18 2015, rana, Update, SUS, Oplevs huh? burp.png
       Reply  Mon Mar 23 13:27:33 2015, ericq, Update, SUS, ITMX oplev quadrant gains unbalanced olsum.png
          Reply  Fri Apr 3 14:57:38 2015, manasa, Update, SUS, BS oplev BS_oplev_Apr3.png
             Reply  Fri Apr 3 15:15:55 2015, Steve, Update, SUS, BS oplev 
                Reply  Fri Apr 3 19:35:14 2015, Jenne, Update, SUS, BS oplev centered 
Message ID: 11125     Entry time: Mon Mar 9 18:10:59 2015     In reply to: 11056
Author: Jenne 
Type: Update 
Category: ASC 
Subject: Oplev filters re-copied 

Back on Feb 20th (elog 11056) Q replaced all of our oplev parts with the aLIGO version. 

Unfortunately, after this it has seemed like there was something not quite right with the optical lever servos.

  • When we would restore kicked optics, after the osem rms values come down the scripts try to engage the optical levers.  This would often kick and ring up the optics (I've seen this with ETMX and ETMY, and once or twice with the beam splitter)  Not good.
  • Sometimes, if the optic wasn't kicked, it would oscillate.  I would see this in particular with ETMY, by seeing the green transmission at the PSL table oscillating.
  • Q noticed that sometimes when the oplevs' outputs were turned off, the outputs were railed at the limiter values.

Since, when the models were changed which gave us an extra underscore in the oplev names, Q did a find-and-replace in the foton text files, I was worried that this might have broken things.  I'm not entirely sure how it would have broken them (I didn't see any difference in a diff), but I've heard enough horror stories about the delicacy of the foton text files.

Anyhow, I opened the last archived foton files from just before Q made the change, and copy-and-pasted the design strings from the old filter banks to the new ones.  Hopefully this fixes things.

ELOG V3.1.3-